OpenStudy (kl0723):

Integral help:

OpenStudy (kl0723):

\[\int\limits x^3+x^4\tan(x)dx\]

OpenStudy (kl0723):

can u guys give me advice about what to pick for u when doing substitution or if I would need to set it up before doing so

OpenStudy (xapproachesinfinity):

first use linearity of integrals: \[\int f(x)+g(x)~dx=\int f(x)dx+\int g(x)dx\]

OpenStudy (kl0723):

I have limits on this problem, could I do this without the limits and then apply the limits of integration after that?

OpenStudy (xapproachesinfinity):

so you have \[\int x^3dx+\int x^4\tan xdx\]

OpenStudy (kainui):

This doesn't look too easy to be honest, without limits it makes it harder, definitely tell us as much info as possible about the problem @kl0723

OpenStudy (kl0723):

ok my limits are from -pi/4 to pi/4

OpenStudy (xapproachesinfinity):

hmm the second part i think you need integration by part as a start

OpenStudy (zarkon):

with those limits the problem is trivial

OpenStudy (kl0723):

according to the book I should be able to apply substitution without the use of "by parts" method, but I don't really know how to set up the problem

OpenStudy (zarkon):

review the integration of odd functions from -a to a

OpenStudy (xapproachesinfinity):

oh i see

OpenStudy (xapproachesinfinity):

i didn't pay attention to odd even thingy hehe

OpenStudy (solomonzelman):

4 times by parts? it's very annoying... teachers!!! ha

OpenStudy (xapproachesinfinity):

just check your textbook you should be able to see it :)

OpenStudy (xapproachesinfinity):

@SolomonZelman there is not by part

OpenStudy (solomonzelman):

I thought I saw that somewhere in the thread.

OpenStudy (zarkon):

if \(f(x)\) is an odd function defined on [-a,a] then \[\int\limits_{-a}^{a}f(x)dx=0\]

OpenStudy (kl0723):

oh wow :P

OpenStudy (xapproachesinfinity):

yes 0 is the solution :)

OpenStudy (kl0723):

@xapproachesinfinity @Kainui and @Zarkon Thank you! I'll make sure to review this

OpenStudy (xapproachesinfinity):

welcome :) i didn't do anything though the credit is to @Zarkon and @Kainui lol

OpenStudy (kl0723):

what if I had an even function?

OpenStudy (zarkon):

if \(f(x)\) is even then \[\int\limits_{-a}^{a}f(x)dx=2\int\limits_{0}^{a}f(x)dx\]

OpenStudy (xapproachesinfinity):

it would be 2int(f(x))dx from [0,a]

OpenStudy (zarkon):

you would still need to do work

OpenStudy (zarkon):

odd are the best ;)

OpenStudy (xapproachesinfinity):

hehe sure they are lol i don't want go integrate by part all that stuff hehe

OpenStudy (zarkon):

in what I wrote above we are obviously assuming the \(f\) is integrable.

OpenStudy (zarkon):

just to be complete

OpenStudy (kainui):

 Here's a picture, just a random odd function and even function. You can see how the area under the curve for an odd function will exactly cancel left and right, while an even function you have double on left and right, so you can just double half the integral, it's not something you should spend time memorizing.